### v1

@@ 4.4 Relation to the erupting filament
To better understand the white-light morphology of the CME and its small-scale structures, we correlated the CME with the solar surface structures and their dynamics, as observed in EUV. This event can be clearly associated with a filament eruption from the southwestern limb of the Sun. Figure 9 shows a polar projection view of the composite SUVI images in the 304Å  and the 171Å  filters. We note that the filament begins crossing the western solar limb on December 4, and we can follow its height over the surface. The eruption and first appearance of the CME in COR1 occurs on December 8, 2022 at $\sim$1 UT. In the lower corona ($<$ 5 ${}_{\odot}$) the filament rises at a speed of 11$\pm$4 km/s, reaching about 310$\pm$18 km/s (kinematics are not shown) after the eruption ($>$ 9 ${}_{\odot}$). As the filament lifts off, a hot (green, 195 Å ) ’horn’-shaped structure (Figure 9, white arrow) offers evidence of a MFR, followed by cool (red, 304Å) radial structures highly reminiscent of filament fine structures.
Figure 9: Polar view of the solar prominence and its eruption in composite images of 304Å  and 171Å  observed by SUVI from December 4 to December 8. In the y-axis the height in solar radii, while in the x-axis the position angle is given. For orientation, in the top left corner a sketch of the PA is given, where 0° corresponds to the Solar North. The image is produced by JHelioviewer (Müller et al., 2017). The animation corresponding to this figure is available online in the file movie2.mov.
Figure 10: Evolution of the CME between the EUV instruments and coronagraphs. Panels (A)-(D) show composite EUVI 171 (blue)/304Å(red) images and STA/COR1 on December 8 between 01:23 and 11:39 UT. The white arrow indicates the hot ’horn’-shape and the red arrow points to the cool filament material. Panels (F) and (G) show the features tracked in WISPR images with COR2 identifications (see Figure 2).
In Figure 10, we show combined EUVI and coronagraph images where we can clearly follow the evolution of these features from EUV to the white-light images. Specifically, panels (A)–(D) show composite EUVI 171/304Å images with COR1 data on December 8 between 01:23 and 11:39 UT. The white arrow indicates the hot ’horn’ shape, while the red arrow points to the cool filament material. In panels (E) and (F) we show COR2 images one hour apart from each other. We mark on those images the WISPR tracked features that can be identified in the COR2 images. Combining EUV and coronagraph data, we associate features 0, 1 and 2 with the typical shapes that we find in filaments during their evolution. Features 3 and 4 instead appear to be related to the core of the CME, indicated by the red arrow in the COR1 images (panels (A)-(D)). In Figure 1 panel (C), we also mark with red arrows thread patterns observed in C3 that might be related to features 3 and 4 as observed in WISPR images (panel (D) from Figure 1, or in more detail shown in Figure 2).
@@ 5 Discussion and conclusions
The internal structure of CMEs has been explored in the past but the new generation of white-light instruments, such as WISPR, deliver images with much higher sensitivity thanks to the unprecedented orbit of PSP. These images hold interesting results, not fully explored yet, about characteristics of the MFR evolving into interplanetary space.
@@
We study the CME event on December 8, 2022 that is related to an eruptive filament. The eruption occurred close to the western limb of the Sun and is observed in white-light by multiple spacecraft located at different distances and vantage points. While the STA and SoHO spacecraft, close to 1 AU and separated by 14°, remain almost constant in their position, PSP is close to the Sun and changes in distance from 0.16 AU to 0.11 AU and in longitude by $\sim$24° during the CME observing time of 32 hours. Using the fast motion of PSP, we investigate WISPR data using a 3D reconstruction method as described in Section 3 and Appendix A. We track 10 well-observed small-scale density structures in WISPR images (Figures 2,  3 and  4). The results enable us to categorize the tracked features and to locate them in the context of the global CME appearance derived from GCS reconstruction.
@@
Based on their morphology, propagation direction, and dynamics we find three different groups of features. Group 1 covers a C- or J-shaped appearance (features 0, 1, 2, 5), group 2 are density patterns characterized by aligned threads (features 3, 4), and group 3 covers blobs (features 6, 7, 8, 9). C- and J-shapes features can be interpreted as plasma-containing twisted magnetic field structures, commonly observed in CMEs. Concave-upward features are frequently observed in EUV, e.g., in our study by SUVI and EUVI (see the ’horns’ pointed with white arrows in Figures 9 and 10), and may be indicators of the extent of the flux rope cavity (Vourlidas et al., 2013). The EUV horns have an appearance similar to the C- and J-shapes observed in WISPR, which suggests coronal temperatures for these features. The high-density blobs observed are usually related to plasmoid instabilities in the post-CME current sheet (see, e.g., Ciaravella & Raymond, 2008; Schanche et al., 2016; Di Matteo et al., 2019; Poirier et al., 2023; Webb & Vourlidas, 2016; Lee et al., 2020). We find numerous blobs propagating in the rear part of the CME under study in a time window of approximately 22 hours. The blobs are derived to move slower by about 100 km/s with respect to the features tracked at the front of the CME.
@@
The thread-like features (group 2) appear localized on a global fishbone structure with a certain angle to that backbone. As these trace the magnetic field fine structure, they might refer to some sheared field with confined density enhancement. They likely resemble cool plasma material as observed in EUV at chromospheric temperatures (see red arrow in Figure 9 and 10), evolving into localized structures during the filament eruption. This can be compared to other prominence eruption studies. For example Su et al. (2015) reports a series of vertical threads aligned along an elongated structure, together with associated U-shape and blob features likely to be related to signatures of magnetic reconnection. In general, we notice that the high-resolution data from WISPR allow to relate in more detail CME small-scale structures to typical elements of a filament (see e.g., Parenti, 2014). This might shed more light on the early evolution of the internal structure of CMEs and contributes to a better understanding of why such a low number of CMEs covering prominence material is detected in-situ (see e.g., Niembro et al., 2023).
@@
Relating the tracked magnetic structures to the global CME structure, we find interesting results. To reconstruct the CME flux rope structure in 3D we used the GCS model (see Figure 8). We obtain that all features of group 1 are clearly located within the CME’s eastern part of the CME. Group 2 is also found to be clearly located in the western leg of the CME. In comparison to that, the blobs from group 3 are distributed over a large longitudinal range of $\sim$40° and do not show a clustered behavior. With respect to the derived GCS geometry, all blobs (features 6–9) are located at the edges of the CME eastern and western flanks, respectively. Features 8 and 9 even outside of the reconstructed flux rope. This may hint toward either the spatial extent of the post-CME current sheet, hence its three-dimensionality, or to the possible influence of the ambient corona in the evolutionary behavior of the CS (e.g., Bemporad et al., 2006; Lin et al., 2015) leading eventually to a secondary source for the formation of the blobs. A clear distinction between these two processes is not possible due to the uncertainties related to both, the PSP reconstruction and GCS method (see Section 3). But we encourage further studies on relating small-scale density structures to the global CME appearance for better constraining their physical generation processes.
@@
The combined views from SoHO, STA, and PSP allow us to compare the 3D shape of the global CME appearance from 1 AU and from 0.11-0.16 AU. Interestingly, we found that WISPR does not detect the outer CME envelope, in contrast to the HI-1 images that clearly show the typical bubble-like front (see Figure 7). Because PSP is much closer to the CME than SoHO and STA, WISPR line-of-sight through the CME is much shorter than for the 1 AU imagers. Therefore, the broad outer front is being ’resolved out’ in WISPR images. On the other hand, due to its closer proximity, WISPR gives higher signals for small-scale structures in the CME body (see also e.g., Vourlidas et al., 2016; Nisticò et al., 2019; Hess et al., 2020, 2021). We also notice that similar effects, namely a different appearance of coronal EUV waves at the solar surface as viewed from different vantage points, were explained by the integration of intensity along different lines-of-sight (e.g., Patsourakos & Vourlidas, 2009).
@@
We conclude that CME flux ropes might consist of different magnetic fine structures that evolve differently. These results are relevant for future CME propagation and forecasting models, especially those that implement magnetized flux ropes (e.g., Török et al., 2018; Verbeke et al., 2019). To accurately interpret the observed density structures of a CME, which trace magnetic structures, it is therefore essential to gather image data from various distances. This approach offers complementary perspectives on the CME’s appearance, accounting for projection and line-of-sight integrated density effects that influence its visual presentation.
@@ Appendix A Single-spacecraft reconstruction fits
As described in Section 3, PSP can be used as a single-spacecraft to obtain the kinematic and position of features belonging to the observed CME. The CME small-scale structures we investigate are faint to be revealed and tracked through the WISPR-I and WISPR-O FoVs, therefore we use so-called “LW-processed” images (described in detail in Appendix A in Howard et al., 2022). This approach exploits the time domain at each pixel location to obtain a smooth background (at the 5th-percentile level) in a time interval centered at the time of each image.
@@
Figure 11: On the top, an example of the grid used by the IDL routine wispr_camera_coords to measure the projected angles in WISPR images, respectively the angle out of the plane $\beta$ and the elongation $\gamma$. On the bottom, a schematic of the geometry used taken from (Liewer et al., 2019), where PSP’s orbit is approximated to be in the solar equatorial plane, and through which we derive the geometrical equations (Equations  3 and 4).
@@
Figure 11 describes the geometry for reconstructing trajectories of small-scale structures in the WISPR FoV. We are interested in their location with respect to the Sun at a specific time t, that is given by $r_{2}(t)$. In addition, we want to obtain for each feature its longitude $\phi_{2}$ and its latitude $\delta_{2}$ in a heliocentric solar equatorial coordinate system, such as the heliocentric inertial reference frame (HCI). Approximating PSP’s orbit to be in the solar equatorial plane, PSP’s coordinates are given by [$r_{\rm PSP},\phi_{\rm PSP},\delta_{\rm PSP}]=[r_{1},\phi_{1},0]$. PSP observes a point-like feature in the WISPR FoV that is identified by two projected angles, the elongation $\gamma$ in PSP’s orbital plane and the angle out of this plane $\beta$, as illustrated in Figure 11. These angles are obtained by tracking the feature over time and converting its coordinate from pixel to PSP orbital frame coordinates using the SolarSoft/IDL routine wispr_camera_coords, accounting for effects of spacecraft location, projection and distortion of the instruments (see more details in Braga & Vourlidas, 2021). Each of the identified features was tracked at least three times, to ensure the track of the same feature over time. This allows to estimate the uncertainties in the results obtained according to the approach followed by Braga & Vourlidas (2021). From these different tracks, Liewer et al. (2020) obtains the average $\beta$ and $\gamma$ at each time, and also estimates the uncertainties in the $\beta$ and $\gamma$ measurements, and uses these uncertainties to estimate the errors in the parameters of the reconstructed trajectory of a feature.
@@
From the measured $\gamma(t)$ and $\beta(t)$ of a specific feature, we determine the unknowns $r_{2}$ (defined as $r_{2}(t)=r_{0}+v_{0}*t$, with $v_{0}$ and $r_{0}$ its constant speed and position at the initial time $t_{0}$), $\phi_{2}$, and $\delta_{2}$ by fitting the following two geometrical equations:
@@
$\beta(t)={\rm atan}\bigg{(}\frac{\tan\delta_{2}\sin(\gamma(t))}{\sin[\phi_{2}-\phi_{1}(t)]}\bigg{)}\leavevmode\nobreak\ ,$ (1)
$\centering\cot(\gamma(t))=\frac{r_{1}(t)-r_{2}(t)\cos\delta_{2}\cos([\phi_{2}-\phi_{1}(t)])}{r_{2}(t)\cos\delta_{2}\sin([\phi_{2}-\phi_{1}(t)])}\leavevmode\nobreak\ ,\@add@centering$ (2)
@@
These two equations provide the conversion between the observed positions ($\beta$ and $\gamma$) of a feature with respect to WISPR and the feature’s 3D position in the heliocentric inertia reference frame, assuming that PSP’s orbit is in the solar equatorial plane. Note that PSP’s orbit is inclined to the solar equator by a small angle ϵitalic-$\epsilon\approx$ 4°. The above equations are therefore modified to include the correction due to the inclination $\epsilon$.
@@
$\beta(t)={\rm atan}\bigg{(}\frac{\tan\delta_{2}\sin(\gamma(t))}{\sin[\phi_{2}-\phi_{1}(t)]}(1-F\sin\epsilon)\bigg{)}\leavevmode\nobreak\ ,$ (3)
@@
where $G$ and $F$ specify the first order corrections, and their expressions are given in the Appendix of Liewer et al. (2020). With PSP’s position ($\phi_{1}$, $r_{1}$ and $\epsilon$) known, and assuming a radial motion at a constant speed, a feature’s trajectory ($\phi_{2},\delta_{2},r_{0}$ and $v_{0}$) can be determined from a sequence of $\gamma(t)$ and $\beta(t)$ measurements.
@@
So far, two different approaches have been adopted by the scientific community to solve these equations in order to obtain its unknowns. Liewer et al. (2020) solves the equations containing the $\epsilon$ term, 3 and 4, performing a Levenberg–Marquardt least-squares fit (see Appendix B of Liewer et al., 2020, for a discussion on deriving initial guess parameters). Instead, Braga & Vourlidas (2021) solves the simplified equations, and finds the best $\phi_{2}(t)$ and $\delta_{2}(t)$ by substituting simultaneously in Equations 1 and 2 the initial guess for each unknown that is set to vary within a certain range that is decided from the user a priori. Then the results are converted to the system of reference of the spacecraft in order to consider the inclination of PSP to the Solar equatorial plane. Once the equations are solved per each set of initial guess, the best fit is computed using the residuals. Indeed, the solution is the one that allows to derive the $\beta_{der}$ and $\gamma_{der}$, through Equations 1 and 2, which have the smallest deviation with respect to the measured ones, $\beta_{mes}$ and $\gamma_{mes}$. In this study, we use both approaches.
@@
Figures 12 and 13 show the related calculation steps and plots to the solving schemes for Equations 3 and 4 using the two different approaches by Liewer et al. (2020) and Braga & Vourlidas (2021), respectively, for all tracked features (0–9). The features were carefully selected considering a significant variation of the longitude of PSP during its observation to satisfy the condition for applying the reconstruction methods (Section 3).
@@
Figure 12 shows for each feature the plots of the evolution of the elongation $\gamma$ and $\beta$ over time. The measured angles and their 1-$\sigma$ uncertainties, calculated considering the three different tracked performed per each substructure, are shown in black. In red we shown the computed $\gamma$ and $\beta$ using the best-fitting parameters derived by forward-fitting Equations 3 and 4. The $\beta$ angle can be also computed with a different equation (see Liewer et al., 2020), which is shown in blue.
Figure 13 shows per each feature to the left the best value of the unknown parameters obtained with the residuals technique. We notice how per each feature the best parameter chosen to perform the fit on the measurement is the one where the minimum of the residual function is located. The right panels show instead the plots of the evolution of the elongation $\gamma$ and $\beta$ over time. The red crosses indicate the measurements performed of the angles, while the solid black line shows the fit.
Figure 12: Results obtained for features (0 to 9) using the approach of (Liewer et al., 2020).
@@
Figure 13: Results obtained for features (0 to 9) using the approach of (Braga & Vourlidas, 2021). The red crosses are the measured angles $\alpha$ and $\beta$, while the solid black line indicated the fit on the equations discussed in the Appendix A with the parameters shown in Table 1.
@@
In this study, the trajectories of the tracked features are determined with two approaches, and in both cases, we assume that a feature moves radially with a constant speed; namely, $\phi_{2}$, $\delta_{2}$, $v_{0}$ are all constant. A relatively large deviation of the computed $\gamma$ and $\beta$ from the measurements, as shown in a few panels in Figure 12 may indicate that the motion of some of these features is not exactly radial at times. We will discuss these deviations in future work. The experiments on deriving time-dependent longitude and latitude are described by (Braga & Vourlidas, 2021).

### v2

@@ 4.4 Relation to the erupting filament
To better understand the white-light morphology of the CME and its small-scale structures, we examined the structure and dynamics of the CME eruption from its coronal origin observed in EUV. This event can be clearly associated with a filament eruption from the southwestern limb of the Sun. Figure 9 shows a polar projection view of the composite SUVI images in the 304Å  and the 171Å  filters. We note that the filament began crossing the western solar limb on December 4, and we can follow its height over the surface. The eruption and first appearance of the CME in COR1 occurs on December 8, 2022, at $\sim$1 UT. In the lower corona ($<$ 5 R), the filament rises at a speed of 11$\pm$4 km/s, reaching about 310$\pm$18 km/s (kinematics are not shown) after the eruption ($>$ 9 R). As the filament lifts off, a hot (green, 195 Å ) ”horn-shaped” structure (Fig.  9, white arrow) offers evidence of an MFR, and it is followed by cool (red, 304Å) radial structures highly reminiscent of filament fine structures.
Figure 9: Polar view of the solar prominence and its eruption in composite images of 304Å  and 171Å  observed by SUVI from December 4 to 8. In the y-axis, the height is in solar radii, while in the x-axis, the position angle is given. For orientation, in the top-left corner, a sketch of the PA is given, where 0° corresponds to the Solar North. The image was produced by JHelioviewer (Müller et al., 2017). The animation corresponding to this figure is available online in the file movie2.mov.
Figure 10: Evolution of the CME between the EUV instruments and coronagraphs. Panels (A)-(D) show composite EUVI 171 (blue)/304Å(red) images and STA/COR1 from December 8 between 01:23 and 11:39 UT. The white arrow indicates the hot horn shape, and the red arrow points to the cool filament material. Panels (F) and (G) show the features tracked in WISPR images with COR2 identifications (see Fig.  2).
In Fig. 10, we show combined EUVI and coronagraph images where one can clearly follow the evolution of these features from the EUV to the white-light images. Specifically, panels (A)–(D) show composite EUVI 171/304Å images with COR1 data on December 8 between 01:23 and 11:39 UT. The white arrow indicates the hot horn shape, while the red arrow points to the cool filament material. In panels (E) and (F), we show COR2 images one hour apart from each other. We mark on those images the WISPR tracked features that can be identified in the COR2 images.
Combining the EUV and coronagraph data, we associated features 0, 1, and 2 with the typical shapes that we are found in filaments during their evolution. Features 3 and 4 instead appear to be related to the core of the CME, indicated by the red arrow in the COR1 images (panels (A)-(D)). In Fig. 1 panel (C), we also mark with red arrows the thread patterns observed in C3 that might be related to features 3 and 4 as observed in WISPR images (panel D from Fig. 1 and shown in more detail in Fig.2).
@@ 5 Discussion and conclusions
The internal structure of CMEs has been explored in the past, but the new generation of white-light instruments, such as WISPR, deliver images with a much higher sensitivity thanks to the unprecedented orbit of PSP. These images hold interesting results, not yet fully explored regarding the characteristics of the MFR evolving into interplanetary space.
@@
We studied the CME event on December 8, 2022, related to an eruptive filament. The eruption occurred close to the western limb of the Sun and was observed in white-light by multiple spacecraft located at different distances and vantage points. While the STA and SOHO spacecraft, close to 1 AU and separated by 14°, remained almost constant in their position, PSP was close to the Sun and changed in distance from 0.16 AU to 0.11 AU and in longitude by $\sim$24° during the CME observing time of 32 hours. Using the fast motion of PSP, we investigated the WISPR data using a 3D reconstruction method, as described in Sect.  3 and Appendix A. We tracked 10 well-observed small-scale density structures in WISPR images (Figs.  23 and  4). The results enabled us to categorize the tracked features and to locate them in the context of the global CME appearance derived from GCS reconstruction.
@@
Based on their morphology, propagation direction, and dynamics we found three different groups of features. Group 1 covers a C- or J-shaped appearance (features 0, 1, 2, 5), group 2 are density patterns characterized by aligned threads (features 3, 4), and group 3 covers blobs (features 6, 7, 8, 9). The C- and J-shaped features can be interpreted as twisted plasma-containing magnetic field structures, which are commonly observed in CMEs. Concave-upward features are frequently observed in EUV. For example, in our study these features were observed by SUVI and EUVI (see the horns indicated with white arrows in Figs.  910), and they may be indicators of the extent of the flux rope cavity (Vourlidas et al., 2013). The EUV horns have an appearance similar to the C- and J-shapes observed in WISPR, which suggests coronal temperatures for these features. The high-density blobs observed are usually related to plasmoid instabilities in the post-CME current sheet (see, e.g., Ciaravella & Raymond, 2008; Schanche et al., 2016; Di Matteo et al., 2019; Poirier et al., 2023; Webb & Vourlidas, 2016; Lee et al., 2020). We found numerous blobs propagating in the rear part of the CME under study in a time window of approximately 22 hours. The blobs were derived to move slower by about 100 km/s with respect to the features tracked at the front of the CME.
@@
The thread-like features (group 2) appear to be localized on a global fish-bone structure with a certain angle to the ”backbone” of that structure. As these features trace the fine structure of the magnetic field, they might refer to some sheared field with a confined density enhancement. They likely resemble cool plasma material as observed in EUV at chromospheric temperatures (see red arrow in Figs.  910), evolving into localized structures during the filament eruption. This can be compared to other prominence eruption studies. For example, Su et al. (2015) reported a series of vertical threads aligned along an elongated structure together with an associated U-shape and blob features likely related to signatures of magnetic reconnection. In general, we note that the high-resolution data from WISPR allow to relate in more detail CME small-scale structures to typical elements of a filament (see e.g., Parenti, 2014). This enhanced level of detail might shed more light on the early evolution of the internal structure of CMEs and contributes to a better understanding of why such a low number of CMEs covering prominence material is detected in situ (see e.g., Niembro et al., 2023).
@@
When relating the tracked magnetic structures to the global CME structure, we found interesting results. To reconstruct the CME flux rope structure in 3D, we used the GCS model (see Fig.  8). We obtained that all features of group 1 are clearly located within the CME’s eastern part. We also found group 2 to be clearly located in the western leg of the CME. In contrast, the blobs from group 3 are distributed over a large longitudinal range of $\sim$40° and do not show a clustered behavior. With respect to the derived GCS geometry, all blobs (features 6–9) are located at the edges of the CME’s eastern and western flanks, respectively. Features 8 and 9 are even outside of the reconstructed flux rope. This may hint toward either the spatial extent of the post-CME current sheet, and hence its three-dimensionality, or to the possible influence of the ambient corona in the evolutionary behavior of the CS (e.g., Bemporad et al., 2006; Lin et al., 2015) eventually leading to a secondary source for the formation of the blobs. A clear distinction between these two processes is not possible due to the uncertainties related to both the PSP reconstruction and the GCS method (see Sect. 3). Nevertheless, we encourage further studies that will relate small-scale density structures to the global CME appearance in order to better constrain their physical generation processes.
@@
The combined views from SOHO, STA, and PSP allowed us to compare the 3D shape of the global CME appearance from 1 AU and from 0.11-0.16 AU. Interestingly, we found that WISPR does not detect the outer CME envelope, in contrast to the HI-1 images that clearly show the typical bubble-like front (see Fig. 7). Because PSP is much closer to the CME than SOHO and STA, the WISPR line of sight through the CME is much shorter than for the 1 AU imagers. Therefore, the broad outer front is ”resolved out” in WISPR images. On the other hand, due to its closer proximity, WISPR gives higher signals for small-scale structures in the CME body (see also e.g., Vourlidas et al., 2016; Nisticò et al., 2019; Hess et al., 2020, 2021). We also note that similar effects, namely a different appearance of coronal EUV waves at the solar surface as viewed from different vantage points, were explained by the integration of intensity along different lines of sight (e.g., Patsourakos & Vourlidas, 2009).
@@
We conclude that CME flux ropes might consist of different magnetic fine structures that evolve differently. These results are relevant for future CME propagation and forecasting models, especially those that implement magnetized flux ropes (e.g., Török et al., 2018; Verbeke et al., 2019). To accurately interpret the observed density structures of a CME, which trace magnetic structures, it is essential to gather image data from various distances. This approach offers complementary perspectives on the CME’s appearance, accounting for projection and line-of-sight integrated density effects that influence its visual presentation.
@@ Appendix A Single-spacecraft reconstruction fits
As described in Sect. 3, PSP on its own can be used to obtain the kinematic information and position of features belonging to an observed CME. The CME’s small-scale structures we investigated are too faint to be revealed and tracked through the WISPR-I and WISPR-O FoVs. Therefore, we use the so-called LW-processed images (described in detail in Appendix A in Howard et al. 2022). This approach exploits the time domain at each pixel location to obtain a smooth background (at the fifth-percentile level) in a time interval centered at the time of each image.
@@
Figure 11: Observational and orbit geometry parameters. Top: Example of the grid used by the IDL routine wispr_camera_coords to measure the projected angles in WISPR images. Respectively, the angle out of the plane is $\beta$ and the elongation is $\gamma$. Bottom: Schematic of the geometry used taken from (Liewer et al. 2019) where PSP’s orbit is approximated to be in the solar equatorial plane and through which we derived the geometrical equations (Eqs. ( 3)- (4)).
@@
Figure 11 describes the geometry for reconstructing trajectories of small-scale structures in the WISPR FoV. We are interested in their location with respect to the Sun at a specific time t, given by $r_{2}(t)$. In addition, we wanted to obtain for each feature its longitude $\phi_{2}$ and its latitude $\delta_{2}$ in a heliocentric solar equatorial coordinate system, such as the HCI reference frame. Approximating PSP’s orbit to be in the solar equatorial plane, its coordinates are given by [$r_{\rm PSP},\phi_{\rm PSP},\delta_{\rm PSP}]=[r_{1},\phi_{1},0]$. PSP observes a point-like feature in the WISPR FoV that is identified by two projected angles: the elongation $\gamma$ in PSP’s orbital plane and the angle out of this plane $\beta$, as illustrated in Fig. 11. These angles were obtained by tracking the feature over time and converting its coordinate from pixel to PSP orbital frame coordinates using the SolarSoft/IDL routine wispr_camera_coords while accounting for effects of spacecraft location, projection, and distortion of the instruments (see more details in Braga & Vourlidas 2021). Each of the identified features was tracked at least three times to ensure the track of the same feature was correctly followed over time. This allowed us to estimate the uncertainties in the results obtained according to the approach followed by Braga & Vourlidas (2021). From these different tracks, Liewer et al. (2020) obtained the average $\beta$ and $\gamma$ at each time, and estimated the uncertainties in the $\beta$ and $\gamma$ measurements, and used these uncertainties to estimate the errors in the parameters of the reconstructed trajectory of a feature.
@@
From the measured $\gamma(t)$ and $\beta(t)$ of a specific feature, we determined the unknowns $r_{2}$, defined as $r_{2}(t)=r_{0}+v_{0}*t$ (with $v_{0}$ and $r_{0}$ as its constant speed and position at the initial time $t_{0}$); $\phi_{2}$; and $\delta_{2}$ by fitting the following two geometrical equations:
@@
$\beta(t)={\rm atan}\bigg{(}\frac{\tan\delta_{2}\sin(\gamma(t))}{\sin[\phi_{2}-\phi_{1}(t)]}\bigg{)}\leavevmode\nobreak$ (1)
$\centering\cot(\gamma(t))=\frac{r_{1}(t)-r_{2}(t)\cos\delta_{2}\cos([\phi_{2}-\phi_{1}(t)])}{r_{2}(t)\cos\delta_{2}\sin([\phi_{2}-\phi_{1}(t)])}\leavevmode\nobreak\ .\@add@centering$ (2)
@@
These two equations provide the conversion between the observed positions ($\beta$ and $\gamma$) of a feature with respect to WISPR and the feature’s 3D position in the heliocentric inertia reference frame, assuming that PSP’s orbit is in the solar equatorial plane. We note that PSP’s orbit is inclined to the solar equator by a small angle, ϵitalic-$\epsilon\approx$ 4°. The above equations were therefore modified to include the correction due to the inclination $\epsilon$:
@@
$\beta(t)={\rm atan}\bigg{(}\frac{\tan\delta_{2}\sin(\gamma(t))}{\sin[\phi_{2}-\phi_{1}(t)]}(1-F\sin\epsilon)\bigg{)}\leavevmode\nobreak$ (3)
@@
where $G$ and $F$ specify the first order corrections, and their expressions are given in the appendix of Liewer et al. (2020). With PSP’s position ($\phi_{1}$, $r_{1}$, and $\epsilon$) known and assuming a radial motion at a constant speed, a feature’s trajectory ($\phi_{2},\delta_{2},r_{0}$, and $v_{0}$) could be determined from a sequence of $\gamma(t)$ and $\beta(t)$ measurements.
@@
So far, two different approaches have been adopted by the scientific community to solve these equations and obtain the unknowns. Liewer et al. (2020) solves the equations containing the $\epsilon$ term, Eqs. (3)- (4), performing a Levenberg–Marquardt least-squares fit (see Appendix B of Liewer et al. 2020, for a discussion on deriving initial guess parameters). Braga & Vourlidas (2021) solves the simplified equations and finds the best $\phi_{2}(t)$ and $\delta_{2}(t)$ by substituting simultaneously in Eqs. (1)- (2) the initial guess for each unknown, which the user sets to vary within a certain range a priori. The results are then converted to the system of reference of the spacecraft in order to consider the inclination of PSP to the solar equatorial plane. Once the equations are solved per each set of initial guesses, the best fit is computed using the residuals. Indeed, the solution is the one that allows the $\beta_{der}$ and $\gamma_{der}$ to be derived through Eqs. (1)-  (2), with the smallest deviations with respect to the measured $\beta_{mes}$ and $\gamma_{mes}$ values. In this study, we used both of the above-described approaches.
@@
Figures  12 and 13 respectively show the calculation steps and plots related to solving Eqs. (3)- (4) using the two different approaches by Liewer et al. (2020) and Braga & Vourlidas (2021) for all tracked features (0–9). The features were carefully selected considering a significant variation of the longitude of PSP during its observation in order to satisfy the condition for applying the reconstruction methods (Sect.  3).
@@
Figure 12 shows for each feature the plots of the evolution of the elongation $\gamma$ and $\beta$ over time. The measured angles and their 1-$\sigma$ uncertainties, which were calculated considering the tracking performed three times per each substructure, are shown in black. In red we show the computed $\gamma$ and $\beta$ using the best-fitting parameters derived by forward fitting Eqs. (3)- (4). The $\beta$ angle can be also computed with a different equation (see Liewer et al. 2020), which is shown in blue.
Figure 13 shows per each feature the evolution of the elongation, $\gamma$, and angle out of the plane, $\beta$. The black crosses indicate the measurements performed on the angles, while the solid red line shows the best fit. The parameters for the fit were chosen by calculating the minimum of the residual function. We clarify that Braga & Vourlidas (2021) performed the fit on each of the tracks and that the best parameters, shown in Table 1, are the average of the results obtained in the total number of tracks. Specifically, Fig.13 only shows the evolution of the angles in one of the three tracks recorded for each feature. In constrast, Liewer et al. (2020) calculated the average for measurements and then performed the fit. This explains why in Fig. 12 the error bars of each measurement are shown.
Figure 12: Results obtained for features (0 to 9) using the approach of (Liewer et al. 2020).
@@
Figure 13: Results obtained for features (0 to 9) using the approach of (Braga & Vourlidas 2021). The black crosses are the measured angles $\gamma$ and $\beta$, while the solid redline indicates the best fit of the equations discussed in Appendix A, with the parameters shown in Table 1.
@@
In this study, the trajectories of the tracked features were determined with two approaches, and in both cases, we assumed that a feature moves radially with a constant speed, namely, $\phi_{2}$, $\delta_{2}$, $v_{0}$ are all constant. A relatively large deviation of the computed $\gamma$ and $\beta$ from the measurements, as shown in a few panels in Fig. 12, may indicate that the motion of some of these features is not exactly radial at times. We will discuss these deviations in future work. The experiments on deriving time-dependent longitude and latitude are described by (Braga & Vourlidas 2021).